OpenStudy (anonymous):

Suppose an equilateral triangle has sides given by 3y - 4 5y - 18 2y + 3 What is the length of each side of the triangle?

hero (hero):

In an equilateral triangle, all sides are equal, so set any two sides equal to each other and solve for y: 3y-4 = 5y - 18 2y = 14 y = 14/2 y = 7 To find the length of each side, substitute y = 7 into each expression and simplify

OpenStudy (anonymous):

thank you!!!